OpenStudy (ineedtheanswersquick):

Solve each equation. #1

OpenStudy (ineedtheanswersquick):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

here

OpenStudy (ineedtheanswersquick):

6 = a/4 + 2 Need to solve for a

OpenStudy (ineedtheanswersquick):

Please show all work.

OpenStudy (ineedtheanswersquick):

Oh wait

OpenStudy (ineedtheanswersquick):

I don't need the work. Please only give me the final answer.

OpenStudy (alexandervonhumboldt2):

\[6=\frac{ a }{ 4 }+2\] well we can't give answers. i'll give you steps. subtract 2 from both sides: 6-2=a/4+2-2 a/4=4 a=4*4=?

OpenStudy (ineedtheanswersquick):

Please, what is the answer? I have to go right now